OpenStudy (anonymous):

Determine the vertex and the axis of symmetry for the function below: y = -x^2 + 2x + 1 a. vertex: (1, 2) axis of symmetry: y = 1 c. vertex: (2, 1) axis of symmetry: y = 1 b. vertex: (-1, 2) axis of symmetry: x = 1 d. vertex: (1, 2) axis of symmetry: x = 1

OpenStudy (zzr0ck3r):

for \(f(x) = ax^2+bx+c\) \(\text{vertex} = (-\frac{b}{2a},f(-\frac{b}{2a}))\) \(\text{axis of symmetry }\rightarrow x = -\frac{b}{2a}\)